# Near-infrared Variability Detected in the Young Star-Forming Dwarf Galaxy SBS 0335-052E

### v2

@@ 1 Introduction
Despite its extremely low stellar mass and metallicity, possible AGN signatures have been reported in SBS 0335-052E. Spatially-resolved near-infrared (NIR) observations with the Hubble Space Telescope (HST) have revealed that, among the SSCs in SBS 0335-052E, the southern SSCs (referred to as SSC 1 and SSC 2 in the literature; 1997ApJ...477..661T) exhibit excess emission at m, corresponding to hot thermal dust continua (2004ApJS..154..211H; 2008AJ....136.1415R; 2008ApJ...678..804E). Its NIR colors in the Wide-field Infrared Survey Explorer (WISE) (3.4 m) and (4.6 m) bands fall in the regime of AGNs ( in Vega magnitudes; 2012ApJ...753...30S). A 9.7 m silicate absorption feature is clearly observed in the Spitzer mid-infrared spectrum (2004ApJS..154..211H), as observed in obscured type 2 AGNs (e.g., 2007ApJ...654L..49S; 2007ApJ...655L..77H; 2015ApJ...803..110H). The equivalent widths of its polycyclic aromatic hydrocarbon (PAH) features are unusually low, placing it in the same parameter space as AGNs rather than typical star-forming regions (2002AJ....124.1995P; 2004ApJS..154..211H; 2005ApJ...633..706W). However, such hot dust emission, often seen in low-mass dwarf galaxies, has generally been attributed to peculiar dust distributions associated with intense star formation rather than AGNs (e.g., 2016ApJ...832..119H; 2025ApJ...979...36S; 2025arXiv251007303R). Observational and theoretical studies suggest that extreme star-forming activity can mimic the hot dust emission typically ascribed to AGNs (2016ApJ...832..119H; 2018ApJ...858...38S; 2025ApJ...979...36S). Accordingly, hot dust emission of SBS 0335-052E has generally been considered to originate from the intense star formation (2002AJ....124.1995P; 2010ApJ...725.1620A; 2014AA...561A..49H; 2016ApJ...832..119H; 2025arXiv251007303R).
@@ 3 Results
Figure 1 shows the WISE W1 and W2 band light curves of SBS 0335-052E corrected for the systematic offsets (Table 1). Figure 1 indicates that the light curves of SBS 0335-052E show clear time variabilities both in the and bands that are significantly brightened by 0.4 and 0.2 magnitudes corresponding to the luminosity variabilities by % and % from Modified Julian date (MJD) 57000 to 59000 at the 20 and 6 significance levels for the and bands, respectively. Moreover, there is a significant dimming in the band from MJD 55600 to 56800 at the level. We also find higher frequency variations at the 2–3 levels at the epoch around MJD 58900 in the band.
@@ 4.2.3 Relation
We can see in Figure 5 that SBS 0335-052E occupies a uniquely distinct position among the known AGN sample. The host stellar mass of SBS 0335-052E is significantly less massive than that of the nearby AGNs and the JWST-discovered broad-line AGNs with host stellar mass estimates. There are no known broad-line AGNs whose properties are equivalent to those of SBS 0335-052E known to date.
Unfortunately, the current uncertainty in of SBS 0335-052E is too large to determine whether its central BH is significantly overmassive relative to the galaxy’s extremely low stellar mass, and thus to place meaningful constraints on seed BH formation scenarios (e.g., see 2020ARA&A..58..257G, and references therein). Future observations – such as high spatial resolution integral field spectroscopy capable of resolving the galaxy’s BH sphere of influence to analyze stellar or gas dynamics (e.g., 2025arXiv250821748J) – will be crucial for tightly constraining in SBS 0335-052E and for investigating possible seed BH formation mechanisms.
@@ 4.3 Relation to JWST-detected High Redshift Broad-line AGNs
The observational property of exhibiting broad H emission while lacking correspondingly strong X-ray emission, as seen in the JWST-detected broad-line AGNs, is similar to the case of SBS 0335-052E (see Section 4.1.2)444After the submission of our paper to arXiv, 2024MNRAS.535..853J point out that the a broad-line LRD GN-280974 at is particularly resemble to SBS 0335-052E in terms of the X-ray and radio weakness, claiming that SBS 0335-052E is the closest and clearest example of low- analog of the JWST-detected broad-line AGNs.. The X-ray weakness of the JWST-detected broad-line AGNs is generally interpreted as the result of Compton-thick (dust-free) gas absorption due to gases with surrounding the AGNs (e.g., 2024ApJ...969L..18A; 2024arXiv241203653I; 2025MNRAS.538.1921M, but see also 2024arXiv240704777K; 2024ApJ...977L..13B for alternative explanations). This interpretation is consistent with the approach we adopt to explain the X-ray weakness of SBS 0335-052E (Section 4.1.2).

### v3

@@ 1 Introduction
Despite its extremely low stellar mass and metallicity, possible AGN signatures have been reported in SBS 0335-052E. Spatially-resolved near-infrared (NIR) observations with the Hubble Space Telescope (HST) have revealed that, among the SSCs in SBS 0335-052E, the southern SSCs (referred to as SSC 1 and SSC 2 in the literature; 1997ApJ...477..661T) exhibit excess emission at m, corresponding to hot thermal dust continua (2004ApJS..154..211H; 2008AJ....136.1415R; 2008ApJ...678..804E). Its NIR colors in the Wide-field Infrared Survey Explorer (WISE) (3.4 m) and (4.6 m) bands fall in the regime of AGNs ( in Vega magnitudes; 2012ApJ...753...30S). A shallow 9.7 m silicate absorption feature is clearly observed in the Spitzer mid-infrared spectrum (2004ApJS..154..211H), as observed in obscured type 2 AGNs (e.g., 2007ApJ...654L..49S; 2007ApJ...655L..77H; 2015ApJ...803..110H). The equivalent widths of its polycyclic aromatic hydrocarbon (PAH) features are unusually low, placing it in the same parameter space as AGNs rather than typical star-forming regions (2002AJ....124.1995P; 2004ApJS..154..211H; 2005ApJ...633..706W). However, such hot dust emission, often seen in low-mass dwarf galaxies, has generally been attributed to peculiar dust distributions associated with intense star formation rather than AGNs (e.g., 2016ApJ...832..119H; 2025ApJ...979...36S; 2025arXiv251007303R). Observational and theoretical studies suggest that extreme star-forming activity can mimic the hot dust emission typically ascribed to AGNs (2016ApJ...832..119H; 2018ApJ...858...38S; 2025ApJ...979...36S). In addition, in extremely low-metallicity environments such as SBS 0335-052E, the reduced silicate abundance can produce AGN-like shallow silicate features in star-forming optically thick dusty clouds, while the low carbon abundance limits the formation of PAHs and consequently suppresses their emission (e.g., 2004ApJS..154..211H). Accordingly, hot dust emission of SBS 0335-052E has generally been considered to originate from the intense star formation (2002AJ....124.1995P; 2004ApJS..154..211H; 2010ApJ...725.1620A; 2014AA...561A..49H; 2016ApJ...832..119H; 2025arXiv251007303R).
@@ 3 Results
Figure 1 shows the WISE and band light curves of SBS 0335-052E corrected for the systematic offsets (Table 1). Figure 1 reveals, for the first time, clear NIR variability in SBS 0335-052E. The band brightened by about 0.4 mag between Modified Julian Date (MJD) 57000 and 59000, corresponding to a luminosity increase of roughly 50%. A similar trend is seen in the band, which brightened by about 0.2 mag (approximately a 20% increase in luminosity) over the same period. These variations are highly significant, with detection levels of in and in . Prior to this large brightening, the band experienced an earlier bright phase followed by a significant dimming from MJD 55600 to 56800, detected at the level. In addition to these multi-year variations, we also detect higher-frequency variability on roughly 1-year timescales, corresponding to fluctuations at the 2–3 level around MJD 58900 in the band.
@@ 4.2.3 Relation
We can see in Figure 5 that SBS 0335-052E occupies the lowest stellar-mass regime among the known AGN samples. The host stellar mass of SBS 0335-052E is significantly less massive than that of the nearby AGNs and the JWST-discovered broad-line AGNs with host stellar mass estimates. There are no known broad-line AGNs whose properties are equivalent to those of SBS 0335-052E known to date.
Considering its together with the host stellar mass, SBS 0335-052E is consistent with, or lies above, the extrapolated local relation. However, the current uncertainty in of SBS 0335-052E is too large to determine whether its central BH is truly overmassive or consistent with the relation, preventing us from placing meaningful constraints on seed BH formation scenarios (e.g., see 2020ARA&A..58..257G, and references therein). Future observations – such as high spatial resolution integral field spectroscopy capable of resolving the galaxy’s BH sphere of influence to analyze stellar or gas dynamics (e.g., 2025arXiv250821748J) – will be crucial for tightly constraining in SBS 0335-052E and for investigating possible seed BH formation mechanisms.
@@ 4.3 Relation to JWST-detected High Redshift Broad-line AGNs
The observational property of exhibiting broad H emission while lacking correspondingly strong X-ray emission, as seen in the JWST-detected broad-line AGNs, is similar to the case of SBS 0335-052E (see Section 4.1.2)444After the submission of our paper to arXiv, 2024MNRAS.535..853J point out that the a broad-line LRD GN-280974 at particularly resembles SBS 0335-052E in terms of the X-ray and radio weakness, claiming that SBS 0335-052E is the closest and clearest example of low- analog of the JWST-detected broad-line AGNs.. The X-ray weakness of the JWST-detected broad-line AGNs is generally interpreted as the result of Compton-thick (dust-free) gas absorption due to gases with surrounding the AGNs (e.g., 2024ApJ...969L..18A; 2024arXiv241203653I; 2025MNRAS.538.1921M, but see also 2024arXiv240704777K; 2024ApJ...977L..13B for alternative explanations). This interpretation is consistent with the approach we adopt to explain the X-ray weakness of SBS 0335-052E (Section 4.1.2).